16. Given a stockindex with a value of $1100, an anticipated dividend of $27, and a risk free rate of 3%, what should be the value of one futures contract on the index?

Please show working and formula in excel!
Answer #1
Solution:
The value of one futures contract on index 1,106
Working Notes:
The value of one futures contract on index
Expecting pay dividend of $27
= stock index value x ( 1+ rate ) - Value of dividend
= 1100 x ( 1+ 0.03) - 27
= 1133 - 27
=1,106
Notes: Formula for stockindex paying dividend
Future value
=S (1 + (r% x T/12)) - compounded valuee of dividends
In Excel format
A B
1 Value
2 S= Stock index Value 1100
3 r% rate of risk free 3%
4 T = time 12
5 compounded value of dividends 27
6
7 The value of one futures contract on index 1106
Formula used in above tab wee get 1106 =(B2*(1+B3*B4/12)-B5)
Please feel free to ask if anything about above solution in comment section of the question.